Cathy has one dime, two quarters, and three nickels. If Cathy randomly selects one coin, what is the probability that she selects a coin worth less than $0.10?

Answer by Boreal(15235)

Less than $0.10 is NOT $0.10 (less than or equal would be)
She has 6 coins, 3 of which are nickels. The other 3 are greater than or equal to $0.10
3/6 = 1/2=50% is the probability.